# On the Accelerated Expansion of the Universe

### v1

@@
Abstract
If we look at from a quantum perspective, the most natural way in which the universe can be created is in entangled pairs whose time flow is oppositely related. This suggests the idea of the creation of a universe-antiuniverse pair. Assuming the validity of this hypothesis, in this paper, we show that the universe expands in an accelerated manner. The same reasoning holds for the anti-universe as well. This idea does not require any form of dark energy as used in the standard cosmological model of $\Lambda$CDM or in the modified theories of gravity.
Keywords:
Dark Energy; Accelerated Expansion; Quantum Focusing; Bekenstein bound; Relative entropy.
@@ 1 Introduction
The study of type Ia supernova in 1998 by two independent projects, the Supernova Cosmology Project and the High-Z Supernova Search Team revealed an astonishing result that the universe is not only expanding but it is expanding at an accelerating rate [1]. The accelerated expansion of the universe is thought to have begun since the universe entered its dark-energy-dominated era roughly 5 billion years ago. The currently accepted theory of gravity, General Relativity, accounts for this accelerated expansion by introducing a small and positive value of the cosmological constant $\Lambda$. Alternative explanations such as quintessence (see[2] for a review) are also hypothesized as the explanation for cosmic acceleration. Some explanations beyond four dimensions also exist such as the Dvali-Gabadadze-Porrati (DGP) model[3] which argues that gravity behaves as $5D$ at large distances while reducing to that of usual $4D$ gravity at short distances. However, the precise nature of dark energy remains unknown. Alternative explanations are promising but have their own shortcomings, as was argued in[4] that the self-accelerating solution of DGP model[5] contains a ghost mode. Very intriguing arguments are presented in [6, 7, 8, 9] that the universe should exist in pairs with their time oppositely related. In this letter, we show using recently conjectured Quantum Focusing[10] that if the universe exists as a pair of universe-antiuniverse pairs, it must expand in an accelerated manner. We begin our discussion by reviewing Quantum Focusing Conjecture (QFC).
@@ 1.1 Quantum Focusing Conjecture
In this section, we review the formulation of Quantum Focusing Conjecture (QFC)
1.1.1 Generalised Entropy for Cauchy-splitting surfaces
@@
Figure 1: A Cauchy surface $\Sigma$ is divided into two parts by a surface $\sigma$. $S_{out}$ is defined as entropy restricted to any one side of splitting surface $\sigma$.
@@ 1.1.2 Quantum Focusing Conjecture
where $R_{ab}$ is the Ricci tensor, $\sigma_{ab}$ is the shear and $k^{a}$ is the (null) tangent vector to the congruence. This gives QFC as:
$0\geq\Theta^{\prime}=\theta^{\prime}+\frac{4G\hbar}{\mathcal{A}}(S^{\prime\prime}_{out}-S^{\prime}_{out}\theta)=-\frac{1}{2}\theta^{2}-\zeta^{2}-8\pi G\langle T_{kk}\rangle+\frac{4G\hbar}{\mathcal{A}}(S^{\prime\prime}_{out}-S^{\prime}_{out}\theta)$ (6)
@@
$\frac{2\pi\mathcal{A}}{\hbar}\langle T_{kk}\rangle-S^{\prime\prime}_{out}\geq 0$ (7)
@@ 1.2 Relative Entropy and Bekenstein Bound
$tr(\rho_{V}\log\rho_{V})-tr(\rho_{V}\log\rho^{0}_{V})$ (10)
This is simply the statement of the positivity of the relative entropy $S(\rho_{V}||\rho^{0}_{V})$ between the state $\rho_{V}$ and the vacuum state $\rho^{0}_{V}$ and thus the bound holds. For an object in a half-space that is far from the circumscribing boundary, (9) reduces to (8). We will use this limit in the next section.
@@ 2 Towards Accelerated Expansion
where $\rho^{(0)}$ is the vacuum state and $\rho$ is some other state. $T_{kk}=T_{ab}k^{a}k^{b}$ where $k^{a}$ and $k^{b}$ are null vectors. (11) essentially tells that the second derivative of the relative entropy with one of the states being the vacuum state is non-negative. This can be easily verified by noticing that the right-hand side of (11) is simply the statement of QNEC(7). The relative entropy in turn can be written as the statement of the Bekenstein Bound as
@@
where $K_{V}=tr(K\rho_{V})-tr(K\rho^{(0)}_{V})$ with $\rho_{V}$ and $\rho_{V}^{(0)}$ being reduced density matrices defined on half-spatial plane $V$ (whose causal development is the Rindler wedge). More precisely, $V$ is the cut of the Cauchy surface $t=0$ defined by $z\geq 0$. $S_{V}=S_{out}-S_{out}^{(0)}$ is the vacuum subtracted entropy, where, $S_{out}=S(\rho_{V})$ and $S_{out}^{(0)}=S(\rho_{V}^{(0)})$. $K$ is the modular Hamiltonian that acts on the Right Rindler wedge. We now apply this result to the universe as a whole. Since the arguments presented hold locally (for a Rindler horizon), the extension to the entire universe requires a global existence of a causal horizon. This is similar to the existence of the Unruh effect for a local Rindler horizon to that of the Hawking effect for a global event horizon of a black hole. This global extension can be achieved by considering a universe-antiuniverse pair in which case the surface at $t=0$ naturally acts as a causal horizon (Fig.2). The universe runs in the future($t>0$) while the anti-universe runs in the past($t<0$). The universe-antiuniverse pair is classically disconnected but quantum-mechanically entangled with the surface at $t=0$ being the entangling surface. In this case, we can define an entanglement entropy $S(\rho_{V})$ and vacuum entanglement entropy $S(\rho_{V}^{(0)})$ where $\rho_{V}$ is the reduced density matrix of the universe defined on the cut of Cauchy surface by $\gamma$ (see Fig.(3)) obtained by tracing out the degrees of freedom of the anti-universe. The modular Hamiltonian $K$ acts on the half-space defined by the region $t>0$ (universe of the anti-universe pair). Thus, the presence of a universe-antiuniverse pair is required to define the entanglement entropies. Moreover, the pair constitutes the whole space with the universe causally disconnected from the anti-universe, lies in the half-space where $K$ is restricted to act. These two key insights allow the global extension of the above results to the entire universe. Note that the entropy $S_{out}$ refers to the entropy on one side of the Cauchy surface but by unitary time evolution is equivalent to states restricted to a null surface $N$[10, 13]. So, consider a Cauchy surface which is split into two parts by a spatial surface $\gamma$ which lies on the null surface $y$ at $t=0$ (see Fig.(3)). Therefore, states on the Cauchy surface are unitarily equivalent to the states restricted to the null surface $y$. In this case, we make use of the fact that the vacuum entanglement entropy on null cuts is stationary ([10, 14, 13, 15]). Therefore, we can now apply (11) and (12) to the universe as a whole. Using (12), (11) becomes ($S^{\prime\prime}_{out}$ cancels on both sides and $S^{{}^{\prime\prime}(0)}_{out}=0$ since the vacuum entropy is stationary)
@@
The affine parameter $\lambda$ can now be taken to be the cosmic time $t$ and the total energy $E$ of the universe is constant. Considering the circumscribing radius of the universe to be $R$ (can be thought of as an imaginary sphere of radius $R$ circumscribing the universe), then, for the half-space, we have (following Casini) $K_{V}\equiv 2\pi ER$. Therefore, we can finally write (14) as
@@
This equation tells that the second derivative of the circumscribing radius is positive, which means that the universe is expanding in an accelerated manner, all by itself! (provided the universe respects NEC) and does not require any hypothetical form of energy such as dark energy (which is the cosmological constant of $\Lambda$CDM). The same reasoning can be applied to the anti-universe by considering it as the half-space defined by the region $t<0$ so that it also expands in an accelerated manner. Therefore, in general, both the universe and anti-universe are expanding which is consistent with[8].
@@ 3 Conclusion
In this letter, we showed that if the universe exists as a universe-antiuniverse pair, then it expands in an accelerated manner. The same reasoning can be extended to the anti-universe if we take it to be half-space defined by region $t<0$. An important condition that is required to hold is the Null Energy Condition(NEC). A familiar statement exists in the study of black holes where the horizon area is a non-decreasing function of time which is the famous Hawking area theorem. It is interesting to note that Hawking’s area theorem also requires the Null Energy Condition (NEC) to hold. This suggests that the study of causal horizons reveals a deeper understanding of the universe. If our approach is correct, then we need no dark energy to explain the present acceleration of the universe and we may need to modify notions of our standard cosmology.
Conflicts of Interest
The author declares no conflicts of interest.
Acknowledgements

### v2

# Grav. Cosmol. No. issue, year

@@
Grav. Cosmol. No. issue, year
On the Accelerated Expansion of the Universe
Naman Kumar1
School of Basic Sciences, Indian Institute of Technology, Bhubaneswar, Odisha, India, 752050
[…]
11footnotetext: e-mail: namankumar5954@gmail.com
@@ 1 Introduction
The study of type Ia supernova in 1998 by two independent projects, the Supernova Cosmology Project and the High-Z Supernova Search Team, revealed an astonishing result that the universe is not only expanding but it is expanding at an accelerating rate [1]. The accelerated expansion of the universe is thought to have begun since the universe entered its dark-energy-dominated era roughly 5 billion years ago. The currently accepted theory of gravity, General Relativity, accounts for this accelerated expansion by introducing a small and positive value of the cosmological constant $\Lambda$. Alternative explanations, such as quintessence (see [2] for a review), are also hypothesized as the explanation for cosmic acceleration. Some explanations beyond four dimensions also exist, such as the Dvali-Gabadadze-Porrati (DGP) model[3], which argues that gravity behaves as 5D at large distances while reducing to that of usual 4D gravity at short distances. However, the precise nature of dark energy remains unknown. Alternative explanations are promising but have their own shortcomings, as was argued in cite 4 that the self-accelerating solution of the DGP model cite 5 contains a ghost mode. Very intriguing arguments are presented in [6, 7, 8, 9] that the universe should exist in pairs with their time oppositely related. In this Letter, we show using recently conjectured Quantum Focusing[10] that if the universe exists as a pair of universe-antiuniverse pairs, it must expand in an accelerated manner. We begin our discussion by reviewing the Quantum Focusing Conjecture (QFC).
@@ 1.1 Quantum Focusing Conjecture
In this section, we review the formulation of the Quantum Focusing Conjecture (QFC)
1.1.1 Generalized Entropy for
Cauchy-splitting surfaces
@@ 1.1.1 Generalized Entropy for
Figure 1: A Cauchy surface $\Sigma$ is divided into two parts by a surface $\sigma$. $S_{out}$ is defined as entropy restricted to any one side of the splitting surface $\sigma$.
@@ 1.1.2 Quantum Focusing Conjecture
where $R_{ab}$ is the Ricci tensor, $\sigma_{ab}$ is the shear, and $k^{a}$ is the (null) tangent vector to the congruence. This gives QFC as:
$\hskip-20.00003pt{}0\geq\Theta^{\prime}=\theta^{\prime}+\frac{4G\hbar}{\mathcal{A}}(S^{\prime\prime}_{out}-S^{\prime}_{out}\theta)$
$\hskip-20.00003pt{}=-\frac{1}{2}\theta^{2}-\zeta^{2}-8\pi G\langle T_{kk}\rangle$
$\hskip-20.00003pt{}+\frac{4G\hbar}{\mathcal{A}}(S^{\prime\prime}_{out}-S^{\prime}_{out}\theta)$ (6)
@@
$\langle T_{kk}\rangle-\frac{\hbar}{2\pi\mathcal{A}}S^{\prime\prime}_{out}\geq 0$ (7)
@@ 1.2 Relative Entropy and Bekenstein Bound
$tr(\rho_{V}\log\rho_{V})-tr(\rho_{V}\log\rho^{0}_{V})\geq 0$ (10)
This is simply the statement of the positivity of the relative entropy $S(\rho_{V}||\rho^{0}_{V})$ between the state $\rho_{V}$ and the vacuum state $\rho^{0}_{V}$, and thus the bound holds. For an object in a half-space that is far from the circumscribing boundary, (9) reduces to (8). We will use this limit in the next section.
@@ 2 Towards Accelerated Expansion
where $\rho^{(0)}$ is the vacuum state and $\rho$ is some other state. $T_{kk}=T_{ab}k^{a}k^{b}$ where $k^{a}$ and $k^{b}$ are null vectors. (11) essentially tells that the second derivative of the relative entropy with one of the states being the vacuum state is non-negative. This can be easily verified by noticing that the right-hand side of (11) is simply the statement of QNEC(7). The relative entropy, in turn, can be written as the statement of the Bekenstein Bound as
@@
where $K_{V}=tr(K\rho_{V})-tr(K\rho^{(0)}_{V})$ with $\rho_{V}$ and $\rho_{V}^{(0)}$ being reduced density matrices defined on half-spatial plane $V$ (whose causal development is the Rindler wedge). More precisely, $V$ is the cut of the Cauchy surface $t=0$ defined by $z\geq 0$. $S_{V}=S_{out}-S_{out}^{(0)}$ is the vacuum subtracted entropy, where, $S_{out}=S(\rho_{V})$ and $S_{out}^{(0)}=S(\rho_{V}^{(0)})$. $K$ is the modular Hamiltonian that acts on the Right Rindler wedge. We now apply this result to the universe as a whole. Since the arguments presented hold locally (for a Rindler horizon), the extension to the entire universe requires a global existence of a causal horizon. This is similar to the existence of the Unruh effect for a local Rindler horizon to that of the Hawking effect for a global event horizon of a black hole. This global extension can be achieved by considering a universe-antiuniverse pair, in which case the surface at $t=0$ naturally acts as a causal horizon (Fig.2). The universe runs in the future($t>0$) while the anti-universe runs in the past($t<0$). The universe-antiuniverse pair is classically disconnected but quantum-mechanically entangle,d with the surface at $t=0$ being the entangling surface. In this case, we can define an entanglement entropy $S(\rho_{V})$ and vacuum entanglement entropy $S(\rho_{V}^{(0)})$ where $\rho_{V}$ is the reduced density matrix of the universe defined on the cut of Cauchy surface by $\gamma$ (see Fig.(3)) obtained by tracing out the degrees of freedom of the anti-universe. The modular Hamiltonian $K$ acts on the half-space defined by the region $t>0$ (universe of the anti-universe pair). Thus, the presence of a universe-antiuniverse pair is required to define the entanglement entropies. Moreover, the pair constitutes the whole space with the universe causally disconnected from the anti-universe, and lies in the half-space where $K$ is restricted to act. These two key insights allow the global extension of the above results to the entire universe. Note that the entropy $S_{out}$ refers to the entropy on one side of the Cauchy surface, but by unitary time evolution is equivalent to states restricted to a null surface NN[10, 13]. So, consider a Cauchy surface which is split into two parts by a spatial surface $\gamma$ which lies on the null surface $y$ at $t=0$ (see Fig.(3)). Therefore, states on the Cauchy surface are unitarily equivalent to the states restricted to the null surface $y$. In this case, we make use of the fact that the vacuum entanglement entropy on null cuts is stationary ([10, 13, 14, 15]). Therefore, we can now apply (11) and (12) to the universe as a whole. Using (12), (11) becomes ($S^{\prime\prime}_{out}$ cancels on both sides and $S^{{}^{\prime\prime}(0)}_{out}=0$ since the vacuum entropy is stationary)
@@
The affine parameter $\lambda$ can now be taken to be the cosmic time $t$, and the total energy $E$ of the universe is constant. Considering the circumscribing radius of the universe to be $R$ with $R=\infty$ (can be thought of as an infinite imaginary sphere of radius $R$ circumscribing the universe). Then, for this region in the half-space, we have (following Casini) $K_{V}\equiv 2\pi ER$. Therefore, we can finally write (14) as
@@
[…]
@@ 3 Conclusion
In this Letter, we showed that if the universe exists as a universe-antiuniverse pair, then it expands in an accelerated manner. The same reasoning can be extended to the anti-universe if we take it to be a half-space defined by the region $t<0$. An important condition that is required to hold is the Null Energy Condition(NEC). A familiar statement exists in the study of black holes, where the horizon area is a non-decreasing function of time, which is the famous Hawking area theorem. It is interesting to note that Hawking’s area theorem also requires the Null Energy Condition (NEC) to hold. This suggests that the study of causal horizons reveals a deeper understanding of the universe. If our approach is correct, then we need no dark energy to explain the present acceleration of the universe, and we may need to modify notions of our standard cosmology.
Acknowledgments